OpenStudy (anonymous):

Solve by using quadratic formula. x^2-4x+2=0

OpenStudy (anonymous):

Medals will be awarded when I return. Thanks for replying.

OpenStudy (anonymous):

\[(-b \pm \sqrt{b^2-4ac})/2a\] a=1,b=4 and c=2 just plug in the values

myininaya (myininaya):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}=\frac{-4 \pm \sqrt{(-4)^2-4(1)(2)}}{2(1)}\]

OpenStudy (anonymous):

it should be noted that you must first imagine the eqaution as ax+bx+c=0 to make sense of the quadratic formula.

myininaya (myininaya):

\[ax^2+bx+c=0\]

OpenStudy (anonymous):

Thank you to all.